OpenStudy (anonymous):

A swimmer travels from the south end to the north end of a 30.0 m pool in 15.0 s and makes the return trip to the starting position in 19.0 s. What is the average velocity for the roundtrip? a. 2.50 m/s to the south c. 2.50 m/s to the north b. 2.27 m/s to the north d. 0.0 m/s

OpenStudy (anonymous):

it's not A

OpenStudy (cwrw238):

average velocity = total distance travelled ------------------ total time taken

OpenStudy (anonymous):

ok uh velocity = distance/time x = 30/15 x = 30/19

OpenStudy (cwrw238):

total distance = 2*30 meters total time = 15 + 19 seconds

OpenStudy (mathmate):

Velocity = (final distance - initial distance)/time =(0-0)/(15+19) =0

OpenStudy (cwrw238):

strictly speaking , in classical mechanics the formula i gave you is for the average SPEED which is independent of the direction, whereas velocity has both speed and direction.

OpenStudy (cwrw238):

mathmate has it

OpenStudy (anonymous):

@mathmate Joe is on a cross-country trip from New York to Los Angeles. He drives for 45.0 min at 70.0 km/h, 10.0 min at 95 km/h, and 30.0 min at 55.0 km/h, and he spends 20.0 min eating lunch and buying gas. What is the average speed for this part of his trip? a. 67.6 km/h c. 54.7 km/h b. 68.4 km/h d. 53.8 km/h

OpenStudy (anonymous):

@patdistheman Two cars drive in the same direction along a straight road, at a constant speed - one at 70 km/hr and the other at 80 km/hr. Assuming that they start at the same point and at the same time, how much sooner would the faster car arrive at a destination 460 km away? a. 49 minutes c. 5.75 hrs b. 6.6 hrs d. 741 minutes

OpenStudy (anonymous):

I do not see how this can be 0. A swimmer travels from the south end to the north end of a 30.0 m pool in 15.0 s and makes the return trip to the starting position in 19.0 s. What is the average velocity for the roundtrip? a. 2.50 m/s to the south c. 2.50 m/s to the north b. 2.27 m/s to the north d. 0.0 m/s

OpenStudy (anonymous):

is it c @patdistheman

OpenStudy (anonymous):

A total distance of 60 meters is traveled in 34 seconds it would seem the average velocity would be (60/34) = 1.765 m/s.

OpenStudy (mathmate):

@mathmate Joe is on a cross-country trip from New York to Los Angeles. He drives for 45.0 min at 70.0 km/h, 10.0 min at 95 km/h, and 30.0 min at 55.0 km/h, and he spends 20.0 min eating lunch and buying gas. What is the average speed for this part of his trip? a. 67.6 km/h c. 54.7 km/h b. 68.4 km/h d. 53.8 km/h What you would do is to tabulate the time and distances.  Average speed = total distance/total time = 95.8333/1.75 hr. = 54.76 km/hr